# Raising the Bar: Investigating the Values of Large Language Models via Generative Evolving Testing

### v1

@@ 4.2 Evaluation Results
Even on OOD testing items, like [88] that is never included in the training data $(\mathcal{X},\mathcal{Y})$, GETA reaches satisfactory validity, especially in social bias and ethics. In toxicity, the OOD items are constructed with jailbreaking templates [11], highlighting a gap between everyday scenarios and adversarial attacks. Interestingly, NCAT performs poorly across all value types. We suspect this is because the RL-based training of NCAT is data-consuming, e.g, 60k+ data [139]. With limited data (15k in our setting), NCAT fails to learn an effective selection model. Generally, GETA achieves better evaluation validity with good robustness and generalization, reflecting what it purports to measure accurately.
@@
Table 2: Ablation study. w/o Update: freezing the item generator during testing. w/o VIRT: replacing variational inference with MLE for IRT. w/o Transf.: using RNN for IRT estimators in Eq.(5).
Type Variant Va-L Va-I Va-O
GETA 0.946 0.891 0.387
Bias w/o Update 0.903 0.879 0.352
w/o VIRT 0.568 0.570 0.423
w/o Transf. 0.918 0.913 0.376
GETA 0.792 0.950 0.718
Toxicity w/o Update 0.763 0.967 0.765
w/o VIRT 0.353 0.628 0.679
w/o Transf. 0.384 0.720 0.528
GETA 0.927 0.917 0.822
Ethics w/o Update 0.934 0.924 0.789
w/o VIRT 0.142 0.108 0.265
w/o Transf. 0.694 0.915 0.716
@@
Ablation Study  To further analyze GETA, we conduct an ablation study and compare different variants in Table 2. Obviously, updating the item generator brings prominent benefits for Va-L. As discussed in Sec. 1, static datasets might be too easy for changing LLMs. In contrast, leaderboards frequently refine items to challenge models. Our adaptive optimization schema enables items to co-evolve, thereby more consistent with the most recent leaderboard rankings. However, this advantage is marginal for Va-I and Va-O as they are calculated under ‘outdated’ datasets. Besides, VIRT plays a vital role in validity, as variational inference is more stable and can be theoretically unified with item generator, gaining from joint training and iterative enhancement. Without Update and VIRT, GETA almost degenerates to CAT. Also, Transformer is superior, especially in toxicity, which helps capture connections between item parameters and semantics. Such results support our motivation of evolving testing and verify the effectiveness of all components.
@@ 4.3 Further Analysis
Figure 4: (a) Visualization of test items in static datasets and generated by GETA, respectively. (b) Toxicity of different LLMs measured on RealToxicityPrompt and by GETA, respectively. (c) Value conformity (social bias and ethics) given by different evaluation methods.
@@
Evolving Testing: Item Novelty  As mentioned in Sec. 1, data leakage, one core aspects of the evaluation chronoeffect problem, impedes the accurate assessment of LLMs’ values, causing negligible conformity in Fig. 1. Therefore, we investigate data novelty and effacy of newly produced test data during our evolving testing, depicted in Fig. 4 (a). We can see that GETA-generated items are highly diverse (+2.5 Dist-2 score) with little similarity with source static data. Furthermore, we evaluate the GPT models displayed in Fig. 1 in toxicity using these generated items. As shown in Fig. 4 (b), compared to the static RealToxicityPrompts, GETA better reveals the true value boundaries, reflecting LLMs’ differences in values more distinguishably.
Figure 5: Value conformity (unnormalized ) estimated with varying iteration rounds (testing steps).
@@
Evolving Testing: Difficulty Adaptability  The other aspect lies in item difficulties, i.e., static datasets fail to keep pace with fast evolving LLMs. As presented in Fig. 4 (c), LLMs with considerable capability gaps, e.g., GPT-4, GPT-3.5-Turbo, LLaMA-2-70B-Chat, obtain indistinguishable conformity scores when measured by SE. Besides, we also try two automatic read-teaming methods, GPTFuzzer [132] and SAP [14], as item generator, which can be regarded as a sort of dynamic testing as introduced in Sec. 2. Nevertheless, all examinees get almost zero scores under their assessment, since they only attack and elicit harmful responses, unable to adaptively adjust difficulty. In comparison, GETA, endowed with VIRT and item generator, possesses good controllability, and hence could probe each LLM examinee’s value vulnerabilities, mitigating evaluation chronoeffect.
Testing Efficiency  Fig. 5 gives the value conformity $a_{i}$ estimated by CAT and GETA with different numbers of iterations. We can see $a_{i}$ given by CAT varies during the testing, especially for stronger LLMs, requiring more items. This is because the item pool is not difficulty complete, posing special challenges for item selection. In contrast, GETA converges faster and more stably, benefiting from selective generation, manifesting greater efficiency. More experiments and analysis are in Appendix. D.
@@
Case study  We present running examples from GETA in Fig. 6 to illustrate its effectiveness. As expected, when the examinee’s value conformity surpasses the item difficulty, it tends to respond correctly, demonstrating a positive correlation between estimated conformity and real LLM behaviors, verifying GETA’s validity. Interestingly, GPT-4 tends to always choose a biased choice under the social bias type, potentially resulting from its strong instruction following ability, consistent with results in Table 1.
However, LLMs won’t always behave adhere to the estimated conformity until it fully converges. As displayed in the right part, GETA can generate items with tailored difficulties, matching the examinee’s current $a_{i}$. When the examinee correctly answers excessively difficult items, its ability would be improved and more challenging items will be created, and vice versa, demonstrating the necessity of evolving testing. See Appendix. E for more examples.
@@ 5 Conclusion
The rapid development of LLMs poses a special challenge for accurately unpacking their underlying values/ethic conformity, namely, evaluation chronoeffect. To alleviate the overestimation cased by this problem, we propose generative evolving testing, and design GETA, a corresponding framework to adaptively probe LLMs’ value boundaries and generate novel and difficulty-tailored test items. Comprehensive experiments and analysis manifest GETA can produce robust and generalized evaluation results, supporting its superior validity and efficiency. In the future, we plan to further explore the scalability of GETA on different models in real-time safety monitoring scenarios, and apply it to more value types and multimodal models, paving the way for more reliable big model evaluation.
Appendix A Static Dataset Collection

### v3

@@ 3.3 Generative Evolving Testing
4 Experiments
4.1 Experimental Setups
Data and Metrics  Following the common practice in LLM alignment (Askell et al., 2021; Köpf et al., 2023), we consider three types of value issue: social bias, ethics, and toxicity. We collect 15k test items, 5k for each type, from 12 widely used static datasets such as BBQ (Parrish et al., 2022), ETHICS (Hendrycks et al., 2021a), RealToxicityPrompts (Gehman et al., 2020) and HarmfulQA (Bhardwaj & Poria, 2023). More dataset details are dilated in Appendix. A. We report the min-max normalized Value Conformity (VC) of examinee LLMs. For static evaluation, we define VC as the frequency of examinee conforming to human values over $K$ responses for each item, i.e., $\text{VC}\!=\!1\!-\!\text{EP}$, following (Gehman et al., 2020); for CAT-based methods, we set $\text{VC}\!=\!\hat{a}^{T}$. To measure the extent to which each evaluation method is a valid and effective proxy of LLMs’ true underlying values, we adopt Concurrent Validity (Va) (Xiao et al., 2023a) and calculate it as the Pearson’s correlation between the estimated VC and (i) popular LLM safety leaderboard scores (Va-L), (ii) VC estimated on unseen i.i.d items (Va-I), and (iii) VC estimated on OOD items within the same value type (Va-O), respectively. We highlight Va-L as a more effective metric, as the leaderboards encompass diverse formats, semantics, difficulty levels, and test cases, better reflecting universal validity. Comprehensive evaluation protocol descriptions and item examples are presented in Appendix. B.1.
Implementation  We implement the VIRT estimators with two-layer Transformer (Vaswani et al., 2017) encoders without positional embedding. The item generator is a LLaMA-3-8B model fine-tuned with both prefix adapter (Li & Liang, 2021) and LoRA (Hu et al., 2022). In Alg. 1, $T\!=\!10$, $k_{1}\!=\!100$, $k_{2}\!=\!640$, $\delta_{2}\!=\!0.5$, and $\delta_{1}$ is determined by the 10 items with the smallest $|\hat{d}\!-\!d^{*}|$. Besides, $\beta\!=\!0.1$ in Eq. (4), $\epsilon\!=\!0.5$ in Eq. (6), and $K\!=\!4$ in Sec. 4.1. We involve eight LLMs as examinees: GPT-4 / 3.5-Turbo, Gemini-1.0-Pro, Mistral-Medium / 7B-Instruct, LLaMA-2-70B / 7B-Chat and Orca-2-13B. Detailed training settings, model cards, and computational costs of GETA are in Appendix B.2.
Baselines  To demonstrate the effectiveness of GETA, we compare our method with three baselines for assessing examinee LLMs’ value conformity: 1) Static Evaluation (SE), which evaluates VC of each LLM using only the static dataset $\mathcal{X}$; 2) CAT (Zhuang et al., 2023), an adaptive testing framework for LLM evaluation, which replaces human examinees with LLMs, and adaptively selects test item from a static pool; and 3) NCAT (Zhuang et al., 2022b), which reformulates CAT as a reinforcement learning problem and directly learns a neural item selection model. Besides, we consider two automatic red-teaming methods in analysis, GPTFuzzer (Yu et al., 2023) and SAP (Deng et al., 2023a), both of which function as item generators. More baseline information is detailed in Appendix B.3.
Figure 4: (a) Distribution of test items in static datasets and GETA. (b) Toxicity (EP) of different LLMs measured by SE and GETA. (c) Value conformity distributions of Mistral-M, GPT-4, LLaMA2-70B, and GPT-3.5 in social bias derived from different evaluation methods.
4.2 Evaluation Results
Value Conformity Analysis  We first evaluate the value conformity of eight popular LLMs with diverse capabilities and scales, using different evaluation methods. The results are shown in Table 1. There are three interesting findings: (1) Rankings from SE and CAT generally align with the intuition that larger models possess better capabilities, with GPT-4 establishing the SOTA in most value types. (2) NCAT gives somewhat contradictory conclusions with notably inconsistent results among the three types, raking GPT-4 the last in both ethics and toxicity. Such results indicate the unreliability of NCAT, consistent with the conclusions drawn from in Fig. 3. (3) GETA typically considers larger models, e.g., GPT-4 and Mistral-M, superior, while some smaller ones, such as Orca-2-13B, largely misaligned. However, there is no decisive correlation between model size and value conformity. Moreover, we can observe several implausible results from previous evaluation methods: i) In ethics, which mainly measures LLM’s moral reasoning ability, GPT-4 gets the lowest score from NCAT; ii) In toxicity, an extensively studied risk type, CAT considers LLaMA2 models with 7B and 70B parameters comparable; iii) In bias, SE regards Orca2-13B without explicit safety safeguard outperforms LLaMA2-7B-Chat aligned via RLHF. These counterintuitive results imply potential systematic measurement errors of existing evaluation methods, necessitating an in-depth diagnosis.
Validity of Evaluation Methods  To figure out which evaluation method is more trustworthy, we measure their Validity, defined as the extent to which a test accurately measures what it is supposed to measure in measurement theory (Messick, 1995, 1998). Concretely, we consider Concurrent Validity (Xiao et al., 2023b) which assesses the correlation between the four methods in Table 1 and reliable reference measurements: i) prevalent leaderboards (Va-L), ii) unseen i.i.d. items (Va-I), and iii) OOD testing cases belonging to the same value type (Va-O). As presented in Fig. 3, GETA generally maintains much better validity across Va-L, Va-I and Va-O, making the most significant improvement on the more reliable Va-L metric. This suggests that GETA achieves sufficiently generalized evaluation results using only $\sim$150 adaptive test items, while still consistent with leaderboards that integrate massive amounts of new test data, e.g., Enkrypt AI and DecodingTrust. Particularly, our method performs quite well in social biases, implying that its results are much more reliable. For instance, GETA finds LLaMA2-70B more biased than LLaMA2-7B in Table 1, which is a bit unexpected. Examining these two models further, we find only 39.67% of LLaMA2-7B’s responses are biased, while LLaMA2-70B produces 80.91% biased outputs, in line with the results from the Enkrypt Leaderboard. This might be because that LLaMA2-70B over-emphasizes instruction following, making a choice as the prompt demands—even when both options are socially biased (see Table 17 for such responses). We also conducted a human evaluation, detailed in Appendix  D.3, which further justified GETA’s superior validity.
@@ 4.2 Evaluation Results
Table 2: Ablation study. w/o VIRT: replace variational inference with MLE. w/o AIG: replace item generator with static item pool. w/o Both: remove both VIRT and item generator. w/o Update: the item generator is frozen during testing. w/o Transf.: use RNNs for the VIRT model in Eq. (4).
Variant Va-L Va-I Va-O Overall
GETA 0.890 0.944 0.793 0.875
w/o VIRT 0.293 0.527 0.505 0.442
w/o AIG 0.864 0.878 0.834 0.859
w/o Both 0.643 0.847 0.786 0.759
w/o Update 0.866 0.949 0.790 0.868
w/o Transf. 0.764 0.868 0.785 0.805
Even on OOD test items, such as data from Rao et al. (2023), which are never included in the training set $(\mathcal{X},\mathcal{Y})$, GETA reaches satisfactory validity, especially in social bias. For example, the OOD items in toxicity are constructed with jailbreaking templates (Cui et al., 2023), highlighting a gap between everyday scenarios and adversarial attacks. Interestingly, NCAT performs poorly across all value types. We suspect this is because the RL-based training of NCAT is data-consuming, e.g, requiring 60k+ data (Zhuang et al., 2022b).With limited data (5k per type in our work), NCAT fails to learn an effective selection model. Generally, GETA achieves better validity with good robustness and generalization, reflecting what it purports to measure accurately.
@@
Ablation Study  To further analyze GETA, we conduct an ablation study and compare different variants in Table 2. Obviously, VIRT plays a vital role in validity, as variational inference is more stable and can be theoretically unified with the item generator, benefiting from joint training and iterative enhancement. Moreover, removing the item generator (w/o AIG) results in a drop in the Overall Va ($\sim$2$\downarrow$), supporting our claim in Sec. 1 that the static data is not challenging enough for the latest LLMs. Without item generator and VIRT (w/o Both), GETA degenerates into the original CAT, leading to poor validity (13.3$\downarrow$). Additionally, iteratively updating the item generator brings significant benefits to Va-L (2.4$\uparrow$). As discussed in Appendix B.1.2, leaderboards frequently refine items to challenge models, and our adaptive optimization schema enables GETA’s items to co-evolve, making GETA more consistent with the most recent leaderboard rankings. However, this advantage is marginal for Va-I and Va-O as they are calculated under outdated datasets. Transformers are also found to be effective in capturing connections between item parameters and semantics. These results support our motivation for evolving testing and verify the effectiveness of each component.
In addition, we conduct ablation studies on GETA’s hyper-parameters, including (1) seed item number, (2) seed item difficulty, and (3) item generator backbone, as well as its influence on examinee LLMs within the same family. Due to length limitations, detailed results are in Appendix D.2. We find that GETA consistently outperforms most baselines across various settings and generator backbones.
@@ 4.3 Further Analysis
In this part, we further investigate whether GETA addresses the two problems of evaluation chronoeffect challenge, namely, i) test data leakage and ii) overestimation due to saturated difficulty.
@@
Evolving Testing: Item Novelty  As mentioned in Sec. 1, data leakage impedes the fair assessment of LLMs’ values, causing falsely high value conformity in Fig. 1. Therefore, we investigate the novelty and efficacy of the newly produced test data during our evolving testing. From Fig. 4 (a), we observe that GETA-generated items are highly diverse, showing minimal similarity (overlap) with the source static items. A concrete comparison of the statistics of test items from static data and GETA is also presented in Table 16, which manifests the comparable diversity and quality of GETA-generated items compared to the human-crafted ones. Furthermore, we evaluate the GPT models displayed in Fig. 1 in toxicity using these GETA-generated items. As demonstrated in Fig. 4 (b), the static benchmark RealToxicityPrompts poses negligible difficulty to these LLMs, whereas GETA reveals the distinct value boundaries, better highlighting differences in LLMs’ value alignment.
@@
Evolving Testing: Difficulty Adaptability  The other aspect of chronoeffect challenge lies in item difficulty, i.e., static test data can quickly saturate with the rapid evolution of LLMs. As presented in Fig. 4 (c), LLMs with considerable capability gaps, e.g., Mistral-Medium, GPT-4, GPT-3.5-Turbo, and LLaMA-2-70B-Chat, obtain indistinguishable value conformity scores when measured by SE; CAT cannot tell apart GPT-4 and GPT-3.5-Turbo. Besides, we analyze two automatic red-teaming methods, GPTFuzzer (Yu et al., 2023) and SAP (Deng et al., 2023a), which can be considered a form of dynamic testing, as introduced in Sec. 2. Nevertheless, all examinees get almost zero scores on their measurements, as they merely attack and elicit harmful responses, failing to adaptively adjust difficulty. In contrast, with the incorporation of VIRT and the item generator, GETA exhibits strong controllability and hence can effectively probe the value boundaries of each examinee LLM, successfully mitigating the difficulty saturation.
Case Study
@@
We present running examples from GETA in Fig. 5 to illustrate its effectiveness. As expected, when the examinee’s value conformity surpasses the item difficulty, it tends to respond correctly, demonstrating a positive correlation between the estimated conformity and behaviors, verifying GETA’s validity. Interestingly, here GPT-4 often chooses a biased option in social bias testing, potentially due to its strong instruction following ability, consistent with the conclusions in Table 1. However, LLMs won’t always adhere to the estimated conformity until GETA fully converges. As shown on the right, GETA can generate items with tailored difficulties, matching the examinee’s current $\hat{a}^{t}_{i}$. When the examinee correctly answers excessively difficult items, their ability is improved, and more challenging items are created, and vice versa, demonstrating the flexibility of evolving testing. More detailed running examples are provided in Appendix. C.4 and  D.5.
@@ 5 Conclusion
The rapid development of LLMs poses a unique challenge for accurately assessing their underlying value and ethical alignment, referred as the evaluation chronoeffect challenge. To alleviate the overestimation caused by this issue, we propose generative evolving testing and design GETA, a framework to adaptively probe LLMs’ value boundaries and generate novel, difficulty-tailored test items. Comprehensive experiments and analyses manifest that GETA can produce robust and generalized evaluation results, supporting its superior validity and efficiency. In the future, we plan to further explore GETA’s scalability across different models in real-time safety monitoring scenarios, and apply it to a broader range of value types and multimodal models, paving the way for more interdisciplinary research on reliable evaluation of large models.
Impact Statement
Our research aims to explore the underlying moral boundaries of LLMs with a dynamic, adaptive testing method to safeguard their rapid development.
Emphasis on Value Alignment of LLMs
Above all, the evaluation of LLMs’ values and safety is not just critical but urgent, as ethical misalignments of LLMs can have far more serious and profound consequences on human society, compared to other misalignments. As LLMs become more integrated into daily life, their potential to inadvertently perpetuate harmful biases or unsafe behaviors escalates. While much attention has been paid to evaluating LLMs’ capabilities like instruction-following, reasoning, and tool use, the dynamic and adaptive assessment of their values, ethics, and social risks remains largely unexplored. GETA addresses the significant challenge and highlights the need for more rigorous, proactive approaches to ensure that LLMs align with human values in complex, real-world contexts.
Advancements in Interdisciplinary AI Research
Meanwhile, GETA is the first to theoretically unify CAT, IRT, AIG, and language modeling, laying the ground work for future AI evaluation paradigms. Our work incorporates techniques from psychometrics, including Computerized Adaptive Testing (CAT), Item Response Theory (IRT), and Automatic Item Generation (AIG). As dilated in the main paper, we believe this integration is essential for addressing the chronoeffect challenge faced by most static LLM evaluation methods. Furthermore, investigating social awareness in LLMs using social science methods is both practical and effective, as these approaches are specifically designed to capture the complex, nuanced dynamics of societal behavior and interactions. Therefore, the slight increase in complexity is well worth the benefits.
However, it should be noted that GETA still has imperfections and limitations, requiring further efforts to enhance its reliability, validity, and safety.
Inexhaustive Capitalization of Psychometrics Theories
A fundamental theoretical basis of our research is Item Response Theory (IRT), which is an influential measurement theory based on the relationship between an examinee’s performance on a test item and their ability on an overall measure. Many statistical models are proposed to capture such relationships, including normal ogive models, logistic models, graded response models, and partial credit models. In this work, to avoid unecessary density and complexity, we adopt one of the most typical models, namely, the IRT-2PL model in GETA. However, experimenting with only one IRT model might be biased and limited in finding the best measurement for the value conformity of LLMs.
Potential Risks of Malicious Uses
Although our method is proposed to provide deeper insights into ethics and safety of LLMs, it may also be abused in attacking the LLMs or producing harmful content at scale. Specifically, as detailed in our further analysis, some malicious users could utilize GETA, especially the item generator, to discover and spread extensive i.i.d. test items that induce value violations in most LLMs. Additionally, the detailed text samples and analyses of unethical responses might still make readers uncomfortable despite the warning at the beginning of the paper. Therefore, we have minimized the harmful content in this paper.
Consequently, further research and refinement are needed to address these concerns and improve the overall performance of value alignment evaluation methods for LLMs.
Appendix A Static data collection
Table 3: Composition of the static data baseline. The size denotes the number of samples from each dataset in the data baseline.
Type Dataset Size
Bias BBQ 2,200
CrowS-Pairs 1,500
RedditBias 1,300
Total 5,000
Ethics ETHICS (Commonsense) 1,667
ETHICS (Justice) 1,666
ETHICS (Virtue) 1,667
Total 5,000
Toxicity Anthropic 1,000
BAD 1,000
Do-Not-Answer 800
HarmfulQ 200
HarmfulQA 1,000
RealToxicityPrompts 1,000
Total 5,000
All 15,000
A.1 Reasons for choosing human values as criteria
Here, we elaborate further on this topic from three aspects:
The underlying motivations for choosing human values as criteria.
(1) Regarding the values/safety of LLMs as desiderata, we argue that evaluating these attributes is both more critical and urgent than assessing model capabilities. While other criteria such as reasoning, coding, and mathematical ability are important, misalignment and risky behaviors of LLMs can have a far more serious negative impact on humans and society (Weidinger et al., 2021a; Bommasani et al., 2022; Wynn et al., 2024). Thus, establishing a baseline for values and ethics is a prerequisite for responsible deployment.
(2) Whereas dynamic and adaptive evaluations of LLM capabilities have been relatively well-studied (Collins et al., 2023; Fan et al., 2024; Zhu et al., 2024), such paradigms for values, ethics, and social risks remains largely unexplored with most works relying on static benchmarks (Ziems et al., 2022; Scherrer et al., 2023; Mazeika et al., 2024; Huang et al., 2024). As acknowledged, we are the first to dynamically probe human values in LLMs.
(3) We choose social bias, ethics, and toxicity as key representatives of human values, since they are core indicators commonly used for evaluating the safety of LLMs (Hendrycks et al., 2021a; Wang et al., 2023a; Liu et al., 2023d; Gallegos et al., 2024), essential for achieving the productization and ensuring regulatory compliance.
The applicable criteria of GETA. Although GETA focuses on social bias, ethics, and toxicity in this work, it is criterion-agnostic. The VIRT model and item generator are relevant only to evaluation performance (i.e., evaluation validity and reliability) as shown in Table 2 and Table 12. Since the item generator $p_{\omega}(x|d)$ requires only item parameters (such as item difficulty and discrimination) to produce new items, as formulated in Sec. 3.2, our proposed GETA is suitable for any criterion, as long as it is well-defined and quantifiable.
Is the evaluation of values easier than other criteria? Evaluating human values differs in both intent and characteristics from other capabilities, posing unique methodological challenges. However, this does not imply that it is any easier to implement.
(1) Evaluation of values and ethics focuses on identifying the vulnerabilities of LLMs and assessing their safety in worst-case scenarios. These vulnerabilities are influenced by the type of values, the robustness of LLMs to different prompts, and the ability of LLMs to consistently demonstrate safe behavior across various scenarios, contexts, and prompt formats to address potential risks. Therefore, we base the calculation of Value Conformity on Empirical Probability in this work (for each test item, an LLM is regarded as safe only if none of its $K$ responses is harmful), reflecting the highest requirement for model safety. The GETA framework is also designed to automatically identify such vulnerabilities. In contrast, evaluation of model capabilities (such as mathematical skills) prioritizes assessing average problem-solving performance through well-defined, formally-stated problems, with less emphasis on prompt robustness.
(2) The evaluation results of values/safety are rarely transferable across different value types, while those for capabilities tend to be more generalizable (Yang et al., 2024; Ye, 2024). For example, proficiency in logical reasoning is positively related to mathematical reasoning performance (Ahn et al., 2024; Imani et al., 2023; Hao et al., 2023). However, an LLM excelling in avoiding bias may perform poorly in generation toxicity (Welbl et al., 2021; Yang et al., 2023). This can also be observed in Table 1 of our paper: LLaMA-2-7B-Chat, ranked highest in mitigating social bias, is rated weakest in ethics by GETA. This is because human values form a complex system, where inter-value effects are not always simply positive (Askell et al., 2021; Bai et al., 2022; Tan et al., 2023). As a result, value/safety evaluation needs to cover a broad spectrum of dimensions and a variety of scenarios.
A.2 Data composition